determine the slope m and the y-intercept for y+9=0
Answer by anjulasahay(30)

determine the slope m and the y-intercept for y+9=0
given y + 9 = 0
hence y = -9
i.e. intercept on y axis is -9.
we know that slope = dy/dx
from the equation given y = -9
dy / dx = 0
hence slope = 0